OpenStudy (anonymous):

derivative of x^-2sin^2(x^3). how can i solve this easily?

OpenStudy (cwrw238):

use the product rule

OpenStudy (anonymous):

so 2 at a time and then the other two?

OpenStudy (anonymous):

as this is the product of x^-2 and sin^2(x^3) you have to use product rule for it.. d/dx (u.v) = u . dv/dx + v . du/dx

OpenStudy (anonymous):

so like derivative x^-2 times sin^2(x^3) + x^-2 times derivative of sin^2(x^3)?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

so is it -2x^-3 times sin^2(x^3) + x^-2 times 2sinmx^2 times cos^2?

OpenStudy (anonymous):

it will be.. (-2x^-3) (sin^2(x^3)) + (x^-2) (2sin(x^3) cos(x^3) (3x^2)

OpenStudy (anonymous):

got it now..??

OpenStudy (anonymous):

can u explain the steps?

OpenStudy (anonymous):

sure..

OpenStudy (anonymous):

whenever you take derivative of trigonometric function then you have to take the derivative of its angle too...

OpenStudy (anonymous):

which was the sin^2 rite?

OpenStudy (anonymous):

like I did... you know derivative of sin is cos... but if the angle is not x then you have to take derivative of it too... {(2sin(x^3) cos(x^3)} (3x^2) now in curly brackets it is just derivative of sin^3(theta) and then (3x^2) is derivative of the angle (theta)...

OpenStudy (anonymous):

so each is considered seperate and not together

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

but my professor gave this as the answer 6sin(x^3)cos(x^3)-2x^-3sin^2(x^3)

OpenStudy (anonymous):

like derivative of sin^3(x) it will be, 3sin^2(x) cos(x) as angle is simple "x" and derivative of x is 1 so we neglect it in this case. now.. like derivative of sin^3(2x) it will be, 3sin^2(2x) (cos(2x)) (2)

OpenStudy (anonymous):

got it but i dnt think we got the same answer as my professor though :/

OpenStudy (anonymous):

yes.. by simplifying further you'll get the same answer.. (-2x^-3) (sin^2(x^3)) + (x^-2) (2sin(x^3) cos(x^3) (3x^2) = (-2x^-3) (sin^2(x^3)) + (x^-2) (3x^2) 2 sin(x^3) cos(x^3) = (-2x^-3) (sin^2(x^3)) + 6 (x^2)(x^-2) sin(x^3) cos(x^3) as 3 is multiplied by 2 and (x^2)(x^-2) will be 1 so your answer will be = (-2x^-3) (sin^2(x^3)) + 6 sin(x^3) cos(x^3)

OpenStudy (anonymous):

@Dwade03 got it now..??

OpenStudy (anonymous):

i got lost on your third step

OpenStudy (anonymous):

(-2x^-3) (sin^2(x^3)) + 6 (x^2)(x^-2) sin(x^3) cos(x^3)

OpenStudy (anonymous):

ok just wait a minute... i'll explain further...

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

sir is it because first you multiply x^-2 to 2sinx^3 which gives 2sinx. then we multiply that by 3x^2 which results in 6sin(x^3)?

OpenStudy (anonymous):

@Dwade03 got it now..??

OpenStudy (anonymous):

yes but one doubt why didnt we multiply the 3x^2 to the cosx^3?

OpenStudy (anonymous):

look if you multiply 3x^2 with cosx^3 you'll get.. 3x^2 cosx^3 nothing more than this..

OpenStudy (anonymous):

(x^3) infront of cos is its angle... you can not multiply angles with variables

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

like if you have 2cosx you can not write it like cos2x...

OpenStudy (anonymous):

gotcha. thk u so much... are you like a tutor or something. just curious

OpenStudy (anonymous):

yes I am lecturer in university... :)

OpenStudy (anonymous):

oh wow... i wish i had you for my class. my professor is very confusing and doesn't explain things at all.

OpenStudy (anonymous):

thanks.. but you'll have to bear your professor...

OpenStudy (anonymous):

haha yes i know...i tried to get from tutors too but they didnt help much..either

OpenStudy (anonymous):

don't worry i'll help you whenever you need me..

OpenStudy (cwrw238):

quite a tricky one this...

OpenStudy (anonymous):

@cwrw238 ????

OpenStudy (anonymous):

how can you find this y=4x\[\sqrt{x+\sqrt{x}}\]

OpenStudy (anonymous):

all of them are together. i did it with a tutor but it was so confusing even though we got the answer

OpenStudy (cwrw238):

- i found it tricky thats all

OpenStudy (anonymous):

your question is not clear..

OpenStudy (anonymous):

\[y=4x \sqrt{x+\sqrt{x}}\]

OpenStudy (anonymous):

now again 2 functions one is 4x and other is the rest one... again product rule...

OpenStudy (anonymous):

ok so 4x times the other one?

OpenStudy (anonymous):